OpenStudy (anonymous):

When mixing cleaners the directions may say mix 5 parts water to 2 part clorox. Write it as 3 different ways (ratio). If i needed 6 gal of

OpenStudy (anonymous):

if i need 6 gal of water, how much clorox will i need?

OpenStudy (shadowfiend):

So, first off, let's look at the three different ways to write it. One ratio is obviously 5 : 2 -- 5 parts water to 2 parts clorox. Are the other two forms ratios or can they be fractions?

OpenStudy (anonymous):

ratios

OpenStudy (anonymous):

ratios

OpenStudy (anonymous):

so it would be 5 to 2 an 5/2?

OpenStudy (shadowfiend):

Right. Also you can flip it around -- 2 : 5, or 2 parts clorox to 5 parts water.

OpenStudy (anonymous):

okay so on to the second part it gives me a hint to use a proportion to figure the problem out

OpenStudy (shadowfiend):

Right.

OpenStudy (shadowfiend):

So basically you can say `if I have 5 parts of water to 2 parts of clorox, and I have 6 gallons of water, how much clorox do I have'. You can set that up by setting the two fractions equal to each other: \[\frac{5}{2} = \frac{6}{x}\] You can flip that so it's a little easier to solve for \(x\) -- just remember to flip both sides! \[\frac{2}{5} = \frac{x}{6}\]